# MetaML-Pro: Cross-Stage Design Flow Automation for Efficient Deep Learning Acceleration

### v1

@@ Abstract.
This paper presents a unified framework for codifying and automating optimization strategies to efficiently deploy deep neural networks (DNNs) on resource-constrained hardware, such as FPGAs, while maintaining high performance, accuracy, and resource efficiency. Deploying DNNs on such platforms involves addressing the significant challenge of balancing performance, resource usage (e.g., DSPs and LUTs), and inference accuracy, which often requires extensive manual effort and domain expertise. Our novel approach addresses two key issues: cross-stage co-optimization and optimization search. By seamlessly integrating programmatic DNN optimization techniques with high-level synthesis (HLS)-based metaprogramming and leveraging advanced design space exploration (DSE) strategies like Bayesian optimization, the framework automates both top-down and bottom-up design flows, reducing the need for manual intervention and domain expertise. The proposed framework introduces customizable optimization, transformation, and control blocks to enhance DNN accelerator performance and resource efficiency. Experimental results demonstrate up to a 92% DSP and 89% LUT usage reduction for select networks, while preserving accuracy, along with a 15.6-fold reduction in optimization time compared to grid search. These results underscore the novelty and potential of the proposed framework for automated, resource-efficient DNN accelerator designs.
@@ 1. Introduction
The field of deep learning has witnessed unprecedented growth in recent years, driven by the increasing demand for efficient and high-performance applications (LeCun et al., 2015). Consequently, FPGA-based deep neural network (DNN) accelerator design and optimization have gained significant attention (Zhang et al., 2015, 2020). The development of an efficient FPGA-based DNN design requires a diverse skill set that combines expertise in machine learning with low-level knowledge of the target hardware architecture (Sze et al., 2017). Optimizing these DNN designs is a complex process, as it involves balancing competing objectives. On one hand, high accuracy during inference is crucial from an application perspective. On the other hand, the design must be optimized for the underlying hardware architecture, meeting power, latency and throughput requirements while fitting into the FPGA device (Duarte et al., 2018; Coelho et al., 2021). Achieving an optimal balance between these conflicting objectives requires careful consideration and effective optimization strategies (Wang et al., 2019).
@@
Existing optimization techniques for DNNs and hardware have a limitation in that they are often handled in separate stages, with information exchanged only in a top-down manner from the application to hardware. Manual intervention is usually required for bottom-up information flow. This information exchange between multiple stages can significantly enhance the optimization process, as application and hardware optimizations can interact and affect each other. For instance, modifying the DNN model’s architecture could impact the FPGA resource utilization, and optimizing the FPGA design could influence the DNN model’s accuracy (Wang et al., 2019). In addition, there has been limited focus on combining optimization strategies targeting different abstraction levels, such as algorithmic level neural networks and High Level Synthesis (HLS) C++, hindering the selection of the most effective combination of optimization techniques for a given problem. Furthermore, little work focuses on supporting comprehensive exploration of such combination with various configurations of involved optimization techniques towards DNN accelerations.
@@
In this paper, we address these two aforementioned technical challenges: C1 Custom Co-Optimization Strategies and C2 Cross-Stage Optimization Search. First, to navigate the vast design space of hardware-software exploration, this work employs a unique co-optimization approach encompassing three key aspects: (1) optimizing the DNN software model, (2) refining the DNN hardware architecture, and (3) exploring the design space. By encoding each optimization independently and combining them as needed, this approach enables the modular reuse and adaptation of optimizations across benchmarks and hardware platforms. Leveraging programmatic manipulation, we facilitate close coordination between DNN model and hardware design optimizations, significantly expanding the range of design choices. Advanced techniques such as Bayesian Optimization streamline the exploration of the large hardware-software design space, ensuring effective and systematic optimization (Reagen et al., 2017; Mehrabi et al., 2020; Tuli et al., 2023).
@@
Second, developing efficient DNN accelerators requires a coherent strategy that automates the selection, combination, and tuning of optimization techniques across multiple abstraction levels (Ney et al., 2021; Abdelfattah et al., 2020; Yang et al., 2019; Zhang et al., 2022). It is challenging to identify the most effective combination, order and tuning of optimization techniques to ensure optimal outcomes (Zhang et al., 2020; Xu et al., 2020). This work tackles these challenges by employing a cross-stage optimization search approach that coordinates the interaction between different stages of the design flow. By leveraging a combination of top-down and bottom-up optimization strategies, the framework dynamically adapts to the specific requirements of the hardware and the application, guiding the optimization process iteratively and identifying the best optimization techniques and their sequences.
@@
This work utilizes a Bayesian guided nested-loop optimization structure involving both local and global phases, combining very distinct optimization techniques and incorporating feedback from the latter stages of the design flow to refine the process iteratively, as shown in Fig. 1. The proposed cross-stage framework operates across various computational spaces, such as software and High-Level Synthesis (HLS), and can potentially target diverse hardware platforms including FPGAs and ASICs. It supports autonomous custom optimization tasks (e.g., OPT-1 to OPT-4 and potentially more), each constrained by user-defined parameters to optimize performance.
@@
  • A novel co-optimization framework for FPGA-based DNN accelerators. This framework seamlessly integrates programmatic DNN optimizations with HLS-based metaprogramming, coordinated by state-of-the-art Design Space Exploration (DSE) strategies. It supports automated design iteration processes through top-down and bottom-up flows, enabling more efficient optimization and accurate performance predictions by refining DSE estimates with post-HLS data (Section 3).
  • A library of reusable optimization, transformation and control tasks designed to be customizable and flexible, and that can be easily integrated into our co-optimization framework. Some of the tasks in our library are specific to certain applications and/or target technologies, while others are agnostic, providing versatility and adaptability to the framework.
  • Enabling rapid development of customized cross-stage design-flows. The framework introduces novel building tasks that automate the integration and iteration of various stages in the design flow. These components allow for the automation of both high-level and low-level design processes, streamlining the development of tailored solutions for different use cases and improving decision-making accuracy in the optimization process.
  • A comprehensive evaluation of the proposed framework using multiple benchmarks and different optimization strategies. This evaluation will provide insights into the effectiveness of the framework and its optimization modules under different scenarios (Section 5).
@@
This paper expands on our previous studies (Que et al., 2023, 2024a, 2024c) which primarily focus on local optimizations. In (Que et al., 2023), we propose MetaML, an optimization framework, that customizes design flow for DNNs with local optimizations. In (Que et al., 2024a), we introduce control capabilities within the optimization framework, while (Que et al., 2023) focuses on Quantization Heuristic Search (QHS) optimization task for compressing DNN accelerators. However, these prior works are limited in their ability to perform global optimization due to their focus on isolated optimization tasks. The inability to account for interdependencies between optimization tasks across stages mean that the overall design might not have achieved the optimal balance between accuracy, performance, and resource efficiency. This work addresses these limitations by incorporating bottom-up feedback and leveraging Bayesian optimization to enable global optimization across the entire design flow. By adjusting optimization tasks at each stage based on performance feedback, our framework ensures that each optimization choice contributes to a more cohesive and globally optimal solution. This extended approach maintains high accuracy while improving resource efficiency, specifically in resource-constrained environments. We plan to open-source MetaML-Pro to the research community, aiming to inspire further research and development in this area.
@@ 2. Related Work
The field of FPGA-based DNN acceleration has rapidly expanded, leading to the development of numerous optimization techniques and tools. Several co-optimization techniques have been proposed that optimize both algorithm and hardware stages for DNNs on FPGAs, as discussed in various papers (Yang et al., 2019; Zhang et al., 2022; Hao and Chen, 2018; Hao et al., 2019b, a, 2020; Jiang et al., 2020; Dong et al., 2021; Hao et al., 2021) and in hardware-aware neural architecture search studies like (Ney et al., 2021; Abdelfattah et al., 2020). These optimization strategies are often coupled with design space exploration, but are typically hardcoded and cannot be easily changed or customized.
@@
To overcome the limitations of existing optimization techniques for DNNs and hardware, our framework derives fully automated high-level optimization flows backed by reusable target-specific and target-agnostic building blocks. Our approach enables automated top-down and bottom-up flows, allowing for the creation of customized cross-stage optimization strategies for various DNN designs. Moreover, our approach integrates new and existing optimization techniques targeting various levels of abstraction, such as neural networks through graph optimizations and HLS C++ through source-to-source optimizations (Vandebon et al., 2020).
@@ 4. Implementation
4.1. Auto-Pruning Algorithm with Binary Search
@@
Starting at 0% pruning rate, the auto-pruning algorithm obtains initial accuracy $Acc_{p0}$ at step 1 (s1). It then uses a binary search approach, increasing or decreasing the pruning rate based on whether the accuracy loss is within a user-defined tolerance ($\leq\alpha_{p}$). The algorithm terminates when the rate difference is below a threshold ($\beta_{p}$). The number of steps is determined by $1+log_{2}(1/\beta_{p})$. Algorithm search steps and direction are shown in Fig. 8. Both the tolerance ($\leq\alpha_{p}$) and threshold ($\beta_{p}$) values are initially set to 2% in this work.
4.2. Quantization Heuristic Search (QHS) algorithm
The optimization task with the proposed QHS algorithm operates at the HLS C++ level, modifying the fixed-point numerical representations directly at C++ source-level. At this level, there is more direct control over hardware optimizations, resulting in fewer unintended side-effects than making changes at a higher level. This task supports auto quantization, which automatically determines the lowest bitwidth for each layer while maintaining a given level of accuracy loss. Formally, the objective of this task is defined as:
@@ 4.2. Quantization Heuristic Search (QHS) algorithm
It performs mixed-precision quantization for DNNs, allowing separate precisions for the weights, biases and results of layers of a model, taking the advantages of the customizability of FPGAs. While this customizability creates a massive search space of quantization configurations, dependencies within the network can be used to reduce the degrees of freedom of the design space. While other efforts target inter-layer mixed-precision quantization algorithms, our meta-programming based approach programmatically groups a number of cascaded layers described in HLS C++ into one virtual layer based on specific assumptions to reduce the complexity and improve efficiency of this optimization task. For example, if the output of a convolutional layer is 8 bits, there is no benefit in having a directly subsequent max-pooling layer with a 16-bit width. Conversely, if the max-pooling layer requires a bit-width of 4, an equivalently accurate and more efficient design would set the bit-width of the preceding convolutional layer to 4 as well. This work automatically compacts a model into a series of virtual layers using these dependencies. These virtual layers are constructed by grouping weight-layers, such as convolutional or FC layers, together with batch-normalization, pooling, and activation functions on their outputs.
@@
4.3. Scaling Algorithm.
To accommodate a large DNN design on an FPGA, we use the SCALING $O$-task that automatically reduces the layer size while tracking the accuracy loss $\alpha_{s}$. The search stops either when the loss exceeds $\alpha_{s}$. If necessary, $\alpha_{s}$ can be adjusted to achieve further size reduction with minimal impact on accuracy.
4.4. Co-Optimization Workflow
As illustrated in Fig. 5, our co-optimization approach is unique in its ability to programmatically encode three types of optimizations. At the top level, a DSE optimization program is employed, executing a comprehensive exploration strategy and overseeing local optimizations at two (or more) distinct levels of abstraction. At the software DNN level, optimizations are executed on the architecture using tools like Tensorflow or PyTorch. At the HLS DNN level, optimizations are conducted by modifying the HLS code automatically using metaprogramming. Optimizing DNN models across multiple abstraction levels, from software to hardware, leverages complementary strategies that address different bottlenecks within the system. Software optimizations prioritize algorithmic efficiency and reducing computational complexity, while hardware optimizations target improvements in throughput, latency, resource utilization and energy efficiency.
Our co-optimization workflow (Fig. 5) begins with the Exploration space. Here, a DSE program utilizes a DNN model, training data, and configuration parameters to conduct the exploration workflow. This process yields a variety of hardware designs with trade-offs in metrics such as performance and accuracy. A controller process orchestrates this exploration by managing communication between the DSE program and local optimization spaces. Prior to exploration, the DSE program sets up and deploys optimization programs within these spaces. Fig. 5 depicts two such spaces: one for software (SW) utilizing TensorFlow and another for HLS employing Vivado HLS. Each space optimizes the model according to selected strategies, such as pruning or quantization.
4.5. Metaprogramming based Design Optimization
@@
A novel aspect of our co-optimization approach, as previously mentioned, is the integration of metaprogramming into hardware design optimization, diverging from traditional reliance solely on hardware templates (see Fig. 6(a)). These templates, crafted by hardware experts, are inherently limited in generating design variants from a predefined set of parameter values, and require expertise for extending or accommodating new DNN architectures.
In contrast, our approach optimizes hardware design architecture by programmatically modifying unoptimized High-Level Synthesis (HLS) C++ code using Python, expanding the range of possibilities beyond parameter values to generate diverse design variants. This manipulation encompasses tasks such as adjusting HLS pragmas, code restructuring for complexity reduction or performance enhancement, data type alterations for quantization support, and utilization of specialized libraries tailored for specific FPGA devices (Vandebon et al., 2020) (see HLS space in Fig. 5).
Figure 6(b) illustrates an HLS metaprogram in pseudocode for clarity. Metaprograms are able to analyze and instrument C++ code, creating hardware design variants to be considered during the Design Space Exploration (DSE) process. Hence, in our context, a metaprogram is a program that takes as input a C++ program and generates a C++ code variant using source-to-source transformations. In this example, the metaprogram creates a variant by adding an unroll pragma to all outermost for-loops in the kernel function of the original code. The key steps of the metaprogram include: (1) constructing an Abstract Syntax Tree (AST) representation of the C++ code; (2) searching for node pairs (loop, function) in the AST that meet specific criteria; and (3) iterating through the identified loops to insert the pragma code before the loop construct. A more detailed explanation about querying and instrumentation capabilities can be found in (Vandebon et al., 2020).

### v2

@@ Abstract.
This paper presents a unified framework for codifying and automating optimization strategies to efficiently deploy deep neural networks (DNNs) on resource-constrained hardware, such as FPGAs, while maintaining high performance, accuracy, and resource efficiency. Deploying DNNs on such platforms involves addressing the significant challenge of balancing performance, resource usage (e.g., DSPs and LUTs), and inference accuracy, which often requires extensive manual effort and domain expertise. Our novel approach addresses two core key issues: (i) encoding custom optimization strategies and (ii) enabling cross-stage optimization search. In particular, our proposed framework seamlessly integrates programmatic DNN optimization techniques with high-level synthesis (HLS)-based metaprogramming, leveraging advanced design space exploration (DSE) strategies like Bayesian optimization to automate both top-down and bottom-up design flows. Hence, we reduce the need for manual intervention and domain expertise. In addition, the framework introduces customizable optimization, transformation, and control blocks to enhance DNN accelerator performance and resource efficiency. Experimental results demonstrate up to a 92% DSP and 89% LUT usage reduction for select networks, while preserving accuracy, along with a 15.6-fold reduction in optimization time compared to grid search. These results highlight the potential for automating the generation of resource-efficient DNN accelerator designs with minimum effort.
@@ 1. Introduction
The field of deep learning has witnessed unprecedented growth in recent years, driven by the increasing demand for efficient and high-performance applications (LeCun et al., 2015). Consequently, FPGA-based deep neural network (DNN) accelerator design and optimization have gained significant attention (Zhang et al., 2015, 2020). The development of an efficient FPGA-based DNN design requires a diverse skill set that combines expertise in machine learning with low-level knowledge of the target hardware architecture (Sze et al., 2017). Optimizing these DNN designs is a complex process, as it involves balancing competing objectives. On one hand, high accuracy during inference is crucial from an application perspective. On the other hand, the design must be optimized for the underlying hardware architecture, meeting power, latency and throughput requirements while fitting into FPGA devices (Duarte et al., 2018; Coelho et al., 2021). Achieving an optimal balance between these conflicting objectives requires careful consideration and effective optimization strategies (Wang et al., 2019).
@@
Existing optimization techniques for DNNs and hardware have a limitation in that they are often handled in separate stages, with information exchanged only in a top-down manner from the application to hardware. Bottom-up information flow, where hardware design insights inform DNN model optimization, typically requires manual intervention. However, this information exchange across multiple stages can significantly enhance the optimization process, as application and hardware optimizations can interact and affect each other. For instance, modifying the DNN model architecture could impact the FPGA resource utilization, and optimizing the FPGA design could influence the DNN model’s accuracy (Wang et al., 2019). In addition, there has been limited focus on combining optimization strategies targeting different abstraction levels, such as algorithmic level neural networks and High Level Synthesis (HLS) C++, hindering the selection of the most effective combination of optimization techniques for a given problem. In summary, few studies haven thoroughly explored the vast configuration space that emerges when combining DNN optimization techniques across both software and hardware domains.
In this paper, we address these aforementioned technical challenges as follows:
@@
  • C1. Custom Co-Optimization Strategies. First,this work employs a unique co-optimization approach that refines both the DNN software model and the hardware architecture. By encoding software and hardware optimizations independently and composing them as needed, this approach enables modular reuse and adaptation of optimizations across different benchmarks and platforms. Through programmatic manipulation, it ensures tight coordination between DNN model and hardware-level optimizations, significantly expanding the design space and enhancing exploration capabilities.
@@
  • C2. Cross-Stage Optimization Search. Second, developing efficient DNN accelerators requires a coherent strategy that automates the selection, combination, and tuning of optimization techniques across multiple abstraction levels (Ney et al., 2021; Abdelfattah et al., 2020; Yang et al., 2019; Zhang et al., 2022). It is challenging to identify the most effective combination, order and tuning of optimization techniques to ensure optimal outcomes (Zhang et al., 2020; Xu et al., 2020). This work addresses these challenges by employing a cross-stage optimization search approach that coordinates the interaction between different stages of the design flow. By leveraging a combination of top-down and bottom-up optimization strategies, the framework dynamically adapts to the specific requirements of the hardware and the application, guiding the optimization process iteratively and identifying the best optimization techniques and their sequences.
@@
Our proposed co-optimization framework, as illustrated in Fig. 1, operates across various computational spaces, such as software and High-Level Synthesis (HLS), and can potentially target diverse hardware platforms including FPGAs and ASICs (see C1). Each computational space represents an autonomous, custom optimisation task (e.g., OPT-1 to OPT-5, and potentially more), governed by tunable parameters. Moreover, this work utilizes a Bayesian (Reagen et al., 2017; Mehrabi et al., 2020; Tuli et al., 2023) guided nested-loop optimization structure involving both local and global phases, combining very distinct optimization tasks and incorporating feedback from the latter stages of the design flow to refine the process iteratively (see C2).
@@
  • A novel co-optimization framework that enables modular and adaptive hardware-software exploration for FPGA-based DNN accelerators. It independently encodes and combines optimizations for efficient design refinement. A cross-stage optimization search automates selection, ordering, and tuning across abstraction levels, dynamically adapting to hardware and application constraints through top-down and bottom-up strategies, ensuring efficient and high-performance designs (Section 3).
  • A library of reusable optimization, transformation, and control tasks designed to be customizable and flexible, and that can be easily integrated into our co-optimization framework. This library includes optimization tasks such as pruning, quantization, and scaling to enhance DNN accelerator performance and resource efficiency. Moreover, we leverage metaprogramming to programmatically analyze and modify HLS C++ designs, enhancing design flexibility, performance tuning, and resource efficiency. In addition, we integrate Bayesian optimization to accelerate design space exploration while maintaining an optimal balance between accuracy and hardware efficiency. Some of the tasks in our library are specific to certain applications and target technologies, while others remain agnostic, providing versatility and adaptability within the framework (Section 4).
  • A comprehensive evaluation of the proposed framework using multiple benchmarks and different optimization strategies. This evaluation provides insights into the effectiveness of the framework and its optimization modules under different scenarios (Section 5).
To the best of our knowledge, this is the first automated framework that combines reusable optimization components with adaptive feedback and global exploration to jointly balance resource efficiency and accuracy in deep learning designs.
@@
This paper expands on our previous studies (Que et al., 2023, 2024a, 2024c) which primarily focus on local optimizations. In (Que et al., 2023), we propose MetaML, an optimization framework, that customizes design flow for DNNs with local optimizations. In (Que et al., 2024a), we introduce control capabilities within the optimization framework, while (Que et al., 2023) focuses on Quantization Heuristic Search (QHS) optimization task for compressing DNN accelerators. However, these prior studies are limited in their ability to perform global optimization due to their focus on isolated optimization tasks. The inability to account for interdependencies between optimization tasks across stages mean that the overall design might not have achieved the optimal balance between accuracy, performance, and hardware resource. This work addresses these limitations by incorporating bottom-up feedback and leveraging Bayesian optimization to enable global optimization across the entire design flow. By adjusting optimization tasks at each stage based on performance feedback, our framework ensures that each optimization choice contributes to a more cohesive and globally optimal solution. This extended approach maintains high accuracy while improving resource efficiency, specifically in resource-constrained environments. We plan to open-source MetaML-Pro to the research community, aiming to inspire further research and development in this area.
@@ 2. Related Work
The field of FPGA-based DNN acceleration has rapidly expanded, leading to the development of numerous optimization techniques and tools. Several co-optimization techniques have been proposed that optimize both algorithm and hardware stages for DNNs on FPGAs, as discussed in various papers (Yang et al., 2019; Zhang et al., 2022; Hao and Chen, 2018; Hao et al., 2019b, a, 2020; Jiang et al., 2020; Dong et al., 2021; Hao et al., 2021) and in hardware-aware neural architecture search studies like (Ney et al., 2021; Abdelfattah et al., 2020; Fan et al., 2022). While hardware-aware neural architecture search (HW-NAS) approaches typically involve a holistic approach where the search algorithm must consider both software and hardware aspects together, often requiring a broader range of expertise within both NAS and hardware, our work modularizes the optimization process, allowing domain experts to focus on their areas of expertise. In addition, this modular design allows for individual optimization tasks to be reused independently in different contexts or different models, while the holistic nature and tight coupling of HW-NAS makes it challenging to reuse parts of the optimization process independently. This work can independently update and refine specific modules without affecting the entire system, resulting in better reusability and adaptability.
@@
To overcome the limitations of existing optimization techniques for DNNs and hardware, our framework provides fully automated high-level optimization flows backed by reusable target-specific and target-agnostic building blocks. More specifically, our approach enables automated top-down and bottom-up flows, allowing for the creation of customized cross-stage optimization strategies for various DNN designs. Moreover, our approach integrates new and existing optimization techniques targeting various levels of abstraction, such as neural networks through graph optimizations and HLS C++ through source-to-source optimizations (Vandebon et al., 2020).
@@ 4. Implementation
4.1. Co-Optimization Workflow
As illustrated in Fig. 4, our co-optimization approach allows the encoding of optimizations across different models of computation - from software to hardware (see C1 in Section 1). At the top level, a DSE optimization program is employed (see C2), executing a comprehensive exploration strategy and overseeing local optimizations at two (or more) distinct levels of abstraction. At the software DNN level, optimizations are executed on the architecture using tools like Tensorflow or PyTorch. At the HLS DNN level, optimizations are conducted by modifying the HLS code automatically using metaprogramming. Optimizing DNN models across multiple abstraction levels, from software to hardware, leverages complementary strategies that address different bottlenecks within the system. Software optimizations prioritize algorithmic efficiency and reducing computational complexity, while hardware optimizations target improvements in throughput, latency, resource utilization and energy efficiency.
Our co-optimization workflow begins with the Exploration space. Here, a DSE program utilizes a DNN model, training data, and configuration parameters to conduct the exploration workflow. This process yields a variety of hardware designs with trade-offs in metrics such as performance and accuracy. A controller process orchestrates this exploration by managing communication between the DSE program and local optimization spaces. Prior to exploration, the DSE program sets up and deploys optimization programs within these spaces. Fig. 4 depicts two such spaces: one for software (SW) utilizing TensorFlow and another for HLS employing Vivado HLS. Each space optimizes the model according to selected strategies, such as pruning or quantization.
4.2. Auto-Pruning with Binary Search
@@
Starting at 0% pruning rate, the auto-pruning algorithm obtains initial accuracy $Acc_{p0}$ at step 1 (s1). It then uses a binary search approach, increasing or decreasing the pruning rate based on whether the accuracy loss is within a user-defined tolerance ($\leq\alpha_{p}$) as shown in Table 1. The algorithm terminates when the rate difference is below a threshold ($\beta_{p}$). The number of steps is determined by $1+log_{2}(1/\beta_{p})$. Algorithm search steps and direction are shown in Fig. 9. Both the tolerance ($\leq\alpha_{p}$) and threshold ($\beta_{p}$) values are initially set to 2% in this work.
4.3. Auto-Scaling
To accommodate large DNN designs on FPGAs, we adopt the SCALING $O$-task that automatically reduces the layer size while monitoring the result accuracy loss, denoted by $\alpha_{s}$. The objective of this $O$-task is formally defined as:
(2) $\underset{}{\text{minimum}}$ $Scaling\_factor$
subject to $Accuracy\_loss(Scaling\_factor)\leq\alpha_{s}$
The auto-scaling process begins with a scaling factor of 1 (i.e., no scaling), getting initial accuracy $Acc_{s0}$ at step 1 (s1). The algorithm then gradually reduces the scaling factor, continuing the search until the accuracy loss exceeds the threshold $\alpha_{s}$. If necessary, $\alpha_{s}$ can be adjusted to achieve further size reduction with minimal impact on accuracy.
4.4. Auto-Quantization with Heuristic Search
In our framework, we introduce quantization task based on the Quantization Heuristic Search (QHS) algorithm (Que et al., 2024c). It automatically determines the lowest bitwidth for each layer while maintaining a given level of accuracy loss. Formally, the objective of this task is defined as:
@@ 4.4. Auto-Quantization with Heuristic Search
It performs mixed-precision quantization for DNNs, allowing separate precisions for the weights, biases and results of layers of a model, taking the advantages of the customizability of FPGAs. While this customizability creates a massive search space of quantization configurations, dependencies within the network can be used to reduce the degrees of freedom of the design space.
@@
4.5. Metaprogramming for HLS Design Optimization
@@
A novel aspect of our co-optimization approach is the integration of metaprogramming into hardware design optimization. In this section, we describe its key motivation and benefits, and how we integrate metaprogramming in our co-optimization framework.
Limitations of template-based approaches. Traditional FPGA-based DNN acceleration frameworks typically rely on manually crafted templates to generate hardware designs. These templates encode expert knowledge, ensuring optimized implementations for a given set of parameters as illustrated in Fig. 6(a). However, they impose inherent limitations in flexibility, adaptability, and scalability. First, templates generate only a limited number of design variants by adjusting predefined parameters, restricting design exploration. They are optimized for specific architectures and are not easily generalizable to new DNN models or hardware platforms.
Extending these templates to support new DNN architectures requires deep domain knowledge in hardware design, making the process resource-intensive. Hardware engineers must manually introduce modifications, which increases development time and effort. Additionally, the modifications are constrained to specific parameterized components, preventing broader architectural transformations that might be necessary for emerging hardware and application demands. As a result, template-based methods struggle to efficiently explore the vast design space available in FPGA-based DNN acceleration.
Benefits of metaprogramming. Our metaprogramming approach addresses these challenges by dynamically analyzing and modifying HLS C++ source code using Python. This technique allows greater flexibility in exploring hardware optimization strategies. One key advantage is the ability to automate the insertion and tuning of HLS pragmas, such as loop unrolling and pipelining, without requiring manual intervention. By adjusting these pragmas dynamically, the hardware implementation can be fine-tuned for improved performance and resource utilization.
Additionally, metaprogramming facilitates code restructuring to improve computational efficiency and reduce design complexity. By programmatically modifying code, redundant computations can be eliminated, and data dependencies can be optimized. Another important capability is automated data type adjustments for quantization, which optimize memory footprint and computational precision by dynamically adapting numerical representations. Unlike static templates, which rely on predefined precision settings, metaprogramming can analyze the computational requirements of different layers and assign appropriate precision dynamically.
Another significant benefit is the incorporation of specialized FPGA libraries (Vandebon et al., 2021), enhancing performance for target platforms. By integrating device-specific optimizations at the source-code level, our approach ensures that the generated hardware designs are optimized for the underlying FPGA architecture. This adaptability makes metaprogramming a powerful tool for hardware-aware DNN optimization.
Example of an HLS metaprogram. Fig. 6(b) illustrates an HLS metaprogram that dynamically modifies a given C++ source file. The process begins with the construction of an Abstract Syntax Tree (AST) representation of the C++ code. The AST provides a structured representation of the source code, allowing precise modifications for any of its constructs. The metaprogram then performs code querying and analysis to identify outermost loop structures within kernel functions. These loops are critical for performance optimization, as their execution directly impacts resource utilization and latency.
Once the relevant loops are identified, the metaprogram instruments the code by inserting pragma directives such as #pragma unroll \$n. These directives control loop unrolling dynamically, enabling optimized parallel execution of computations. By automating this process, the metaprogram allows rapid exploration of different optimization strategies without requiring manual code modifications. This automation significantly improves both efficiency and design flexibility, making it easier to adapt to different DNN architectures and FPGA platforms. A more detailed explanation about querying and instrumentation capabilities can be found in  (Vandebon et al., 2020).
Figure 7. Integration of the Quantization Heuristic Search (QHS) algorithm with HLS4ML and metaprogramming. The metaprogram dynamically adjusts data types and bit-widths at the HLS C++ level, optimizing computational efficiency while maintaining model accuracy.
Application in quantization optimization. We integrate Quantization Heuristic Search (QHS), detailed in Section 4.4, as an optimization task within our framework using metaprogramming to refine fixed-point numerical representations directly at the HLS C++ level. As illustrated in Fig. 7, we develop a metaprogram called H4ML-API, which programmatically analyzes and modifies HLS4ML-generated code to automate the quantization process.
The workflow proceeds as follows. First, the HLS4ML tool generates an initial, unoptimized firmware from a DNN model. Then, 1 the H4ML-API metaprogram inspects the source structure and executes the firmware to extract DNN architecture details, layer characteristics, and accuracy metrics. This information is 2 passed to the QHS process, which identifies opportunities for bit-width reduction while maintaining model accuracy, dynamically adjusting bit-widths based on layer-specific requirements. The refined bit-width settings are then 3 transferred back to the H4ML-API, which 4 modifies the HLS C++ code accordingly. This iterative process continues from step 1 while the accuracy loss remains within the $\alpha_{q}$ threshold.
A key feature of this approach is the creation of virtual layers to optimize bit-width allocation. Rather than treating each operation separately, the metaprogram groups cascaded operations into logical units, enabling more effective quantization. For instance, if a convolutional layer produces an 8-bit output, subsequent operations such as max-pooling and batch normalization are assigned the same precision, avoiding redundant higher-bit computations. By systematically reducing bit-widths without compromising accuracy, this approach significantly decreases resource usage while maintaining high inference performance.